# A Deep Reinforcement Learning-Based TCP Congestion Control Algorithm: Design, Simulation, and Evaluation

### v1

@@ Abstract
Index Terms:
TCP Congestion Control, Reinforcement Learning, Deep Q-Network, NS-3, Network Simulation, AI in Networking
@@ II Background and Related Work
Network Topology: Network topology refers to the physical or logical arrangements that define the interconnection of devices in a computer network and the logic of data exchange between them. Gains in reliability, performance, energy consumption, and complexity are achieved through network topology [1].
@@
TCP: The purpose of TCP is to ensure secure data transfers in computer networks. Transmitted data reaches the destination completely, in order, and accurately. Data transfer first establishes a connection between the sender and receiver and manages it through a ”three-way handshake” process. If errors occur in the data during transmission, TCP detects and corrects the errors, and retransmits the data. Large data groups are divided into smaller groups and transmitted to the receiver as combined groups [2].
@@
Congestion Window: Network congestion control has evolved significantly since the late 1980s. Jacobson [3] introduced the concept of ”packet conservation” and proposed solutions like slow start and congestion windows to prevent congestion-induced collapses. These innovations were implemented to improve TCP’s performance under congested conditions. Subsequent research has focused on specific performance issues. Cardwell et al. [4] presented BBR as a solution to buffer bloat and low throughput problems associated with loss-based congestion control in networks with large or small bottleneck buffers. The field has expanded to include wireless networks, with various TCP versions and protocols specifically developed for congestion avoidance and control in wireless environments [5]. This evolution reflects the increasing complexity and diversity of modern network infrastructures, indicating a shift in congestion control research from addressing network stability to overcoming specific performance challenges in different network types.
@@
Network Simulator 3: With the growth of computer networks and the development of complex scenarios, network simulators have begun to play a significant role in the research area. Simulators are useful tools for implementing new security solutions, managing time and resources, and evaluating performance. By enabling the virtualization of the environment, they allow the analysis of network performance under different scenarios and their modeling according to specific criteria. NS3 offers features such as working compatibility, robust memory management, debugging, coding in C++, object-oriented concepts, and models supported by NS-2. It is highly suitable for wireless networks [6].
@@
Deep Q-Learning: Deep Q-Learning (DQL) is a significant algorithm in reinforcement learning that uses Deep Q-Networks (DQN) to approximate the Q-function in sequential decision-making problems [7]. Although successful in laboratory conditions, its real-world application is limited due to gaps between theory and practice. DQL has been applied to various fields such as natural language processing [8], image processing, and financial forecasting [9]. The algorithm’s success has led to increased interest in understanding human learning processes and has advantages over other learning algorithms as it does not require prior data [7]. Recent advancements in GPU technology have expanded the application areas of deep learning methods, facilitating the transition from shallow to deep networks [10]. Researchers continue to develop optimized versions of DQL, such as updating neural network weights at each step to improve agent performance during training [7].
@@
Reinforcement Learning: Reinforcement Learning (RL) is a branch of artificial intelligence that focuses on training agents to make decisions by interacting with environments to maximize cumulative rewards [11]. RL consists of fundamental components such as states, actions, policies, and reward signals [11]. Algorithms in RL are classified as model-based, model-free, value-based, and policy-based [11]. Sutton and Barto’s [12] seminal book provides a comprehensive guide covering the history, fundamental concepts, and algorithms of RL. The book consists of three parts: problem definition, basic solution methods, and advanced topics [12]. Recent developments in RL include the rise of deep reinforcement learning, which enhances RL’s capacity to solve complex, high-dimensional problems [11]. RL has found applications in various fields such as autonomous systems, healthcare, and energy management [11].
@@
SOLID Principles: SOLID principles are an acronym combining the Single Responsibility Principle (SRP), Open-Closed Principle (OCP), Liskov Substitution Principle (LSP), Interface Segregation Principle (ISP), and Dependency Inversion Principle (DIP). The main objective of these principles is to increase software quality, reusability, scalability, and maintainability [13, 14].
@@ II-A Network Topologies
An incorrect network topology can lead to problems such as performance degradation, data loss, high energy consumption, and critical system failures. Therefore, selecting the correct network topology is crucial [1]. Network topologies are generally divided into two types:
@@
Star Topology: In this network topology, all devices are arranged around a central hub. Each device connects with others through this central device. The central device regulates and controls data flow. It is frequently preferred due to its simplicity and ease of fault detection, but if the central device fails, the entire system goes down [15]. Dumbbell Topology: In this network topology, there are two ends, and these ends are connected to each other via a main link. Data between these ends is carried over the main link. The ends can behave like a local area network. It is used for traffic analysis and performance testing [2]. Mesh Topology: In this network topology, devices can establish multiple connections with each other. If every device is connected to every other device, it is called a full mesh; otherwise, it is called a partial mesh. It offers high-level data security and backup possibilities. Network resilience is high. Its disadvantages include high cost and complexity [15].
@@ II-B TCP Protocol and Congestion Control
Developed in the 1970s, TCP/IP forms the foundation of modern internet communication [16]. TCP establishes virtual circuits between computers to manage data transmission as byte streams, while IP performs packet routing [17]. The protocol suite covers core network, transport, routing, and application layers [18]. TCP ensures reliable communication by segmenting large data groups, numbering packets, and using acknowledgment mechanisms to prevent data loss [17]. In 1983, TCP/IP replaced the Network Control Protocol as ARPAnet’s official data transport method [19]. Despite its utility, TCP/IP has security vulnerabilities such as IP spoofing and denial-of-service attacks [16]. To mitigate these risks, security enhancements like IPSec, VPNs, and encryption techniques have been developed [16, 19]. TCP/IP continues to evolve with innovations such as IP mobility, multicast, and IPv6, increasing its capacity and importance in modern networks [19]. The congestion window (cWnd) in TCP is an algorithm created to manage network congestion. This window detects the maximum number of packets a TCP flow can have in the network at any given time, directly affecting data transfer rate and network stability [20]. This concept, developed to prevent congestion collapse, ensures network functionality, especially in dense network traffic. Congestion collapse occurs when excessive data traffic overloads network resources, leading to a severe drop in performance [21]. TCP typically interprets packet loss as a congestion-related problem; however, this can cause issues in wireless networks where losses occur due to interference, signal attenuation, or mobility. To overcome such difficulties, modern congestion control algorithms include mechanisms to adaptively adjust the congestion window (e.g., reducing window size during idle or limited sending periods). The initial size and evolution of the congestion window have been re-developed over the years to improve performance and efficiency. In 2010, Allman made specific recommendations for determining and evolving the initial congestion window size, emphasizing a balance between fast data transfer and network stability [22]. Furthermore, innovative methods in congestion control, such as mechanisms like Explicit Congestion Notification (ECN) and hybrid approaches combining loss-based and delay-based control, optimize performance in different network environments [23]. TCP congestion control algorithms are evolving to address new challenges in modern networks. In 5G networks, high user density, large traffic flows, and mmWave communications pose unique obstacles for TCP performance [24]. Traditional loss-based congestion control algorithms struggle with rapid channel oscillations in 5G, leading to buffer bloat and delays [25]. To overcome these issues, researchers are developing dynamic congestion control mechanisms that adapt to network conditions. For example, D-TCP uses an Adaptive Increase/Adaptive Decrease approach based on available bandwidth [26], while FB-TCP uses fuzzy logic to adjust sending rates in urban 5G deployments [27]. These new algorithms aim to improve throughput, reduce packet loss and latency, and enhance fairness in various network environments. Machine learning techniques are also being investigated to optimize TCP congestion control in 5G networks [24]. The evolution of these methods over time ensures TCP remains compatible and effective in the constantly changing network landscape.
@@ II-C Network Congestion and Performance Issues
Network congestion occurs when a network receives more data than it can process, leading to packet loss and unreliable connections. Congestion can result from hardware factors like aging routers, software issues like inefficient protocols, and other factors such as poor network design. Congestion causes a decrease in network performance, an increase in transmission delays, and a reduction in data transfer speed. As networks evolve to support multimedia applications, congestion continues to be a significant problem. Research has focused on understanding the causes of congestion and developing control solutions. Recent studies investigate neural networks for congestion control in various network environments. A deep reinforcement learning approach in wireless networks has shown significant improvements in throughput compared to traditional protocols [28]. For content-centric networks, a neural network-based algorithm has demonstrated an 85.53% increase in throughput by predicting and preventing queue overflow [29]. In wired networks, neural networks show promise in managing heterogeneous traffic flows and improving quality of service [30]. Additionally, neural networks have been applied to automatically adapt new users to existing quality of service policies in backbone networks and enhance traffic management efficiency [31]. These studies collectively demonstrate the potential of neural networks to address congestion control challenges in various network types, offering improved performance and adaptability compared to traditional methods.
@@ II-D Deep Q-Learning
Deep Q-Learning (DQL) combines Q-learning with deep neural networks to address complex reinforcement learning problems with large state spaces [32]. While DQL has shown remarkable success, especially in Atari games [33], it suffers from overestimation of action values in certain situations [34, 33]. To mitigate this issue, researchers have proposed various solutions, including Double Q-learning, which has demonstrated improved performance in various games [33]. Furthermore, dropout techniques have been investigated as a way to reduce variance and overestimation in DQL, showing significant improvements in stability and performance [35]. Despite these advancements, DQL still faces challenges in real-world applications due to factors such as high-dimensional sensory inputs and the need for efficient environmental representations [32]. Ongoing research aims to increase the applicability and effectiveness of DQL in solving complex reinforcement learning problems across various domains.
@@ II-E Reinforcement Learning
Reinforcement Learning (RL) is a branch of artificial intelligence focused on training agents to make decisions by interacting with environments to maximize cumulative rewards [11]. RL consists of fundamental components such as states, actions, policies, and reward signals [11]. Algorithms in RL are categorized as model-based, model-free, value-based, and policy-based [11]. The groundbreaking book by Sutton and Barto [12] provides a comprehensive guide covering the history, fundamental concepts, and algorithms of RL. The book comprises three sections: problem definition, basic solution methods, and advanced topics [12]. Recent developments in RL include the rise of deep reinforcement learning, which enhances RL’s capacity to solve complex, high-dimensional problems [11]. RL has found applications in various fields such as autonomous systems, healthcare, and energy management [11].
@@ II-F1 Classical TCP Algorithms (Reno, CUBIC, BBR)
Reno: Initially operates in ”Slow Start” mode. In this mode, the congestion window (cwnd) typically starts from a small value and then increases exponentially. When network congestion occurs, this growth rate slows down; the exponential increase becomes linear. In cases of packet loss, the lost packet is retransmitted, and the algorithm transitions to fast recovery mode. Reno’s congestion control was first defined by Jacobson, significantly impacting the algorithm’s performance [3, 36]. Limitations and Disadvantages: It always attributes packet loss to network congestion, which is not always a correct assumption. This can lead to Reno unnecessarily slowing down the network and causing performance degradation. Packet loss can also originate from physical network conditions (signal weakness or interference) [37, 38]. In high-latency networks, it cannot react fast enough due to the delay in slow start processes. This typically weakens network performance due to congestion and delay issues observed in remote networks [38]. It is weak in managing multiple packet losses. Reno is designed for single packet losses and experiences performance degradation with multiple losses [38]. In wireless networks, it may misinterpret packet losses caused by physical conditions, further slowing down the network [37].
@@
CUBIC: Initially, the cwnd value increases up to its maximum value. As it approaches, growth slows down, and then increases faster once it has passed. In case of congestion, the window size is reduced, and after congestion clears, it attempts to return to the maximum value [39]. Disadvantages and Limitations: In low-latency networks, it can cause congestion by excessively increasing the window size [39]. It detects network congestion only through packet loss and is insensitive to other causes. For example, it cannot detect differences in network conditions due to bandwidth limitations or changes in network topology [39]. It has a complex calculation mechanism, which creates an additional system resource load. This can negatively impact efficiency, especially on devices with low processing power [40].
@@
BBR: Initially increases the data transfer rate to measure bandwidth and RTT values. To reduce the excess load generated during this phase, the cwnd value is temporarily decreased. The network’s bandwidth is continuously measured, and the location of the bottleneck is identified. To measure and update the RTT value, the cwnd value is periodically decreased, thus preventing buffer bloat. BBR is considered a significant step towards more efficient traffic control [4]. Disadvantages and Limitations: Due to its continuous bandwidth measurement, it can be misleading in situations where conditions change constantly. Especially when there are sudden fluctuations in the network, the BBR algorithm may not yield accurate results [4]. Incorrect detection of bottleneck points can lead to unnecessary reduction in data transfer rate or congestion. This can cause unnecessary data slowdowns in the network [4].

### v2

@@ II Background and Related Work
Network Topology: Network topology refers to the physical or logical arrangements that define the interconnection of devices in a computer network and the logic of data exchange between them. Gains in reliability, performance, energy consumption, and complexity are achieved through network topology [Lee2008].
@@
TCP: The purpose of TCP is to ensure secure data transfers in computer networks. Transmitted data reaches the destination completely, in order, and accurately. Data transfer first establishes a connection between the sender and receiver and manages it through a ”three-way handshake” process. If errors occur in the data during transmission, TCP detects and corrects the errors, and retransmits the data. Large data groups are divided into smaller groups and transmitted to the receiver as combined groups [Subramanya2014].
@@
Congestion Window: Network congestion control has evolved significantly since the late 1980s. Jacobson [Jacobson1988] introduced the concept of ”packet conservation” and proposed solutions like slow start and congestion windows to prevent congestion-induced collapses. These innovations were implemented to improve TCP’s performance under congested conditions. Subsequent research has focused on specific performance issues. Cardwell et al. [Cardwell2017] presented BBR as a solution to buffer bloat and low throughput problems associated with loss-based congestion control in networks with large or small bottleneck buffers. The field has expanded to include wireless networks, with various TCP versions and protocols specifically developed for congestion avoidance and control in wireless environments [Simsek2008]. This evolution reflects the increasing complexity and diversity of modern network infrastructures, indicating a shift in congestion control research from addressing network stability to overcoming specific performance challenges in different network types.
@@
Network Simulator 3: With the growth of computer networks and the development of complex scenarios, network simulators have begun to play a significant role in the research area. Simulators are useful tools for implementing new security solutions, managing time and resources, and evaluating performance. By enabling the virtualization of the environment, they allow the analysis of network performance under different scenarios and their modeling according to specific criteria. NS3 offers features such as working compatibility, robust memory management, debugging, coding in C++, object-oriented concepts, and models supported by NS-2. It is highly suitable for wireless networks [Mohta2023].
@@
Deep Q-Learning: Deep Q-Learning (DQL) is a significant algorithm in reinforcement learning that uses Deep Q-Networks (DQN) to approximate the Q-function in sequential decision-making problems [Tan2020]. Although successful in laboratory conditions, its real-world application is limited due to gaps between theory and practice. DQL has been applied to various fields such as natural language processing [Kucuk2018], image processing, and financial forecasting [Dogan2019]. The algorithm’s success has led to increased interest in understanding human learning processes and has advantages over other learning algorithms as it does not require prior data [Tan2020]. Recent advancements in GPU technology have expanded the application areas of deep learning methods, facilitating the transition from shallow to deep networks [Seker2017]. Researchers continue to develop optimized versions of DQL, such as updating neural network weights at each step to improve agent performance during training [Tan2020].
@@
Reinforcement Learning: Reinforcement Learning (RL) is a branch of artificial intelligence that focuses on training agents to make decisions by interacting with environments to maximize cumulative rewards [Ghasemi2024]. RL consists of fundamental components such as states, actions, policies, and reward signals [Ghasemi2024]. Algorithms in RL are classified as model-based, model-free, value-based, and policy-based [Ghasemi2024]. Sutton and Barto’s [Sutton1998] seminal book provides a comprehensive guide covering the history, fundamental concepts, and algorithms of RL. The book consists of three parts: problem definition, basic solution methods, and advanced topics [Sutton1998]. Recent developments in RL include the rise of deep reinforcement learning, which enhances RL’s capacity to solve complex, high-dimensional problems [Ghasemi2024]. RL has found applications in various fields such as autonomous systems, healthcare, and energy management [Ghasemi2024].
@@
SOLID Principles: SOLID principles are an acronym combining the Single Responsibility Principle (SRP), Open-Closed Principle (OCP), Liskov Substitution Principle (LSP), Interface Segregation Principle (ISP), and Dependency Inversion Principle (DIP). The main objective of these principles is to increase software quality, reusability, scalability, and maintainability [Nesteruk2020, Singh2015].
@@ II-A Network Topologies
An incorrect network topology can lead to problems such as performance degradation, data loss, high energy consumption, and critical system failures. Therefore, selecting the correct network topology is crucial [Lee2008]. Network topologies are generally divided into two types:
@@
Star Topology: In this network topology, all devices are arranged around a central hub. Each device connects with others through this central device. The central device regulates and controls data flow. It is frequently preferred due to its simplicity and ease of fault detection, but if the central device fails, the entire system goes down [Jiang2015]. Dumbbell Topology: In this network topology, there are two ends, and these ends are connected to each other via a main link. Data between these ends is carried over the main link. The ends can behave like a local area network. It is used for traffic analysis and performance testing [Subramanya2014]. Mesh Topology: In this network topology, devices can establish multiple connections with each other. If every device is connected to every other device, it is called a full mesh; otherwise, it is called a partial mesh. It offers high-level data security and backup possibilities. Network resilience is high. Its disadvantages include high cost and complexity [Jiang2015].
@@ II-B TCP Protocol and Congestion Control
Developed in the 1970s, TCP/IP forms the foundation of modern internet communication [Gangane2015]. TCP establishes virtual circuits between computers to manage data transmission as byte streams, while IP performs packet routing [Ros2015]. The protocol suite covers core network, transport, routing, and application layers [Pujolle1997]. TCP ensures reliable communication by segmenting large data groups, numbering packets, and using acknowledgment mechanisms to prevent data loss [Ros2015]. In 1983, TCP/IP replaced the Network Control Protocol as ARPAnet’s official data transport method [Murhammer1999]. Despite its utility, TCP/IP has security vulnerabilities such as IP spoofing and denial-of-service attacks [Gangane2015]. To mitigate these risks, security enhancements like IPSec, VPNs, and encryption techniques have been developed [Gangane2015, Murhammer1999]. TCP/IP continues to evolve with innovations such as IP mobility, multicast, and IPv6, increasing its capacity and importance in modern networks [Murhammer1999]. The congestion window (cWnd) in TCP is an algorithm created to manage network congestion. This window detects the maximum number of packets a TCP flow can have in the network at any given time, directly affecting data transfer rate and network stability [Handley2000]. This concept, developed to prevent congestion collapse, ensures network functionality, especially in dense network traffic. Congestion collapse occurs when excessive data traffic overloads network resources, leading to a severe drop in performance [Welzl2005]. TCP typically interprets packet loss as a congestion-related problem; however, this can cause issues in wireless networks where losses occur due to interference, signal attenuation, or mobility. To overcome such difficulties, modern congestion control algorithms include mechanisms to adaptively adjust the congestion window (e.g., reducing window size during idle or limited sending periods). The initial size and evolution of the congestion window have been re-developed over the years to improve performance and efficiency. In 2010, Allman made specific recommendations for determining and evolving the initial congestion window size, emphasizing a balance between fast data transfer and network stability [Allman2010]. Furthermore, innovative methods in congestion control, such as mechanisms like Explicit Congestion Notification (ECN) and hybrid approaches combining loss-based and delay-based control, optimize performance in different network environments [Kuzmanovic2003]. TCP congestion control algorithms are evolving to address new challenges in modern networks. In 5G networks, high user density, large traffic flows, and mmWave communications pose unique obstacles for TCP performance [Lorincz2021]. Traditional loss-based congestion control algorithms struggle with rapid channel oscillations in 5G, leading to buffer bloat and delays [Abdullah2023]. To overcome these issues, researchers are developing dynamic congestion control mechanisms that adapt to network conditions. For example, D-TCP uses an Adaptive Increase/Adaptive Decrease approach based on available bandwidth [Kanagarathinam2018], while FB-TCP uses fuzzy logic to adjust sending rates in urban 5G deployments [Poorzare2021]. These new algorithms aim to improve throughput, reduce packet loss and latency, and enhance fairness in various network environments. Machine learning techniques are also being investigated to optimize TCP congestion control in 5G networks [Lorincz2021]. The evolution of these methods over time ensures TCP remains compatible and effective in the constantly changing network landscape.
@@ II-C Network Congestion and Performance Issues
Network congestion occurs when a network receives more data than it can process, leading to packet loss and unreliable connections. Congestion can result from hardware factors like aging routers, software issues like inefficient protocols, and other factors such as poor network design. Congestion causes a decrease in network performance, an increase in transmission delays, and a reduction in data transfer speed. As networks evolve to support multimedia applications, congestion continues to be a significant problem. Research has focused on understanding the causes of congestion and developing control solutions. Recent studies investigate neural networks for congestion control in various network environments. A deep reinforcement learning approach in wireless networks has shown significant improvements in throughput compared to traditional protocols [Midhula2024]. For content-centric networks, a neural network-based algorithm has demonstrated an 85.53% increase in throughput by predicting and preventing queue overflow [Bazmi2014]. In wired networks, neural networks show promise in managing heterogeneous traffic flows and improving quality of service [Kumar2013]. Additionally, neural networks have been applied to automatically adapt new users to existing quality of service policies in backbone networks and enhance traffic management efficiency [Topalova2018]. These studies collectively demonstrate the potential of neural networks to address congestion control challenges in various network types, offering improved performance and adaptability compared to traditional methods.
@@ II-D Deep Q-Learning
Deep Q-Learning (DQL) combines Q-learning with deep neural networks to address complex reinforcement learning problems with large state spaces [Awoga2021]. While DQL has shown remarkable success, especially in Atari games [Hasselt2015], it suffers from overestimation of action values in certain situations [Rastogi2020, Hasselt2015]. To mitigate this issue, researchers have proposed various solutions, including Double Q-learning, which has demonstrated improved performance in various games [Hasselt2015]. Furthermore, dropout techniques have been investigated as a way to reduce variance and overestimation in DQL, showing significant improvements in stability and performance [Sabry2019]. Despite these advancements, DQL still faces challenges in real-world applications due to factors such as high-dimensional sensory inputs and the need for efficient environmental representations [Awoga2021]. Ongoing research aims to increase the applicability and effectiveness of DQL in solving complex reinforcement learning problems across various domains.
@@ II-E Reinforcement Learning
Reinforcement Learning (RL) is a branch of artificial intelligence focused on training agents to make decisions by interacting with environments to maximize cumulative rewards [Ghasemi2024]. RL consists of fundamental components such as states, actions, policies, and reward signals [Ghasemi2024]. Algorithms in RL are categorized as model-based, model-free, value-based, and policy-based [Ghasemi2024]. The groundbreaking book by Sutton and Barto [Sutton1998] provides a comprehensive guide covering the history, fundamental concepts, and algorithms of RL. The book comprises three sections: problem definition, basic solution methods, and advanced topics [Sutton1998]. Recent developments in RL include the rise of deep reinforcement learning, which enhances RL’s capacity to solve complex, high-dimensional problems [Ghasemi2024]. RL has found applications in various fields such as autonomous systems, healthcare, and energy management [Ghasemi2024].
@@ II-F1 Classical TCP Algorithms (Reno, CUBIC, BBR)
Reno: Initially operates in ”Slow Start” mode. In this mode, the congestion window (cwnd) typically starts from a small value and then increases exponentially. When network congestion occurs, this growth rate slows down; the exponential increase becomes linear. In cases of packet loss, the lost packet is retransmitted, and the algorithm transitions to fast recovery mode. Reno’s congestion control was first defined by Jacobson, significantly impacting the algorithm’s performance [Jacobson1988, Stevens1994]. Limitations and Disadvantages: It always attributes packet loss to network congestion, which is not always a correct assumption. This can lead to Reno unnecessarily slowing down the network and causing performance degradation. Packet loss can also originate from physical network conditions (signal weakness or interference) [Floyd2004, Padhye1998]. In high-latency networks, it cannot react fast enough due to the delay in slow start processes. This typically weakens network performance due to congestion and delay issues observed in remote networks [Padhye1998]. It is weak in managing multiple packet losses. Reno is designed for single packet losses and experiences performance degradation with multiple losses [Padhye1998]. In wireless networks, it may misinterpret packet losses caused by physical conditions, further slowing down the network [Floyd2004].
@@
CUBIC: Initially, the cwnd value increases up to its maximum value. As it approaches, growth slows down, and then increases faster once it has passed. In case of congestion, the window size is reduced, and after congestion clears, it attempts to return to the maximum value [Ha2008]. Disadvantages and Limitations: In low-latency networks, it can cause congestion by excessively increasing the window size [Ha2008]. It detects network congestion only through packet loss and is insensitive to other causes. For example, it cannot detect differences in network conditions due to bandwidth limitations or changes in network topology [Ha2008]. It has a complex calculation mechanism, which creates an additional system resource load. This can negatively impact efficiency, especially on devices with low processing power [Kurose2021].
@@
BBR: Initially increases the data transfer rate to measure bandwidth and RTT values. To reduce the excess load generated during this phase, the cwnd value is temporarily decreased. The network’s bandwidth is continuously measured, and the location of the bottleneck is identified. To measure and update the RTT value, the cwnd value is periodically decreased, thus preventing buffer bloat. BBR is considered a significant step towards more efficient traffic control [Cardwell2017]. Disadvantages and Limitations: Due to its continuous bandwidth measurement, it can be misleading in situations where conditions change constantly. Especially when there are sudden fluctuations in the network, the BBR algorithm may not yield accurate results [Cardwell2017]. Incorrect detection of bottleneck points can lead to unnecessary reduction in data transfer rate or congestion. This can cause unnecessary data slowdowns in the network [Cardwell2017].